The perimeter of a rectangle is 66. the length is twice the width. What is the length of the diagonl?
Answer by checkley77(12844)

L=2w
P=2L+2W
66=2*2W+2W
66=4W+2W
6W=66
W=66/6
W=11 Aans.
L=2*11=22 ans.
Proof:
66=2*11+2*22
66=22+44
66=66